Find the length and width of a rectangle if its length is thrice its width and its diagonal is 7.
Answer by nerdybill(7384)

Find the length and width of a rectangle if its length is thrice its width and its diagonal is 7.
.
Let w = width
then
3w = length
.
applying Pythagorean theorem:
w^2 + (3w)^2 = 7^2
w^2 + 9w^2 = 49
10w^2 = 49
w^2 = 4.9
w = 2.2 (width)
.
length:
3w = 3(2.21) = 6.6